OpenStudy (republican31):

Solve the question in the comments because that's where I can type equations. How to?

OpenStudy (mathmale):

Please type in the problem you wish to solve.

OpenStudy (republican31):

\[\sqrt{4y+1}-\sqrt{y-2}=3\]

OpenStudy (mathmale):

You'll need to solve for y, but before you attempt that you'll need to square both sides of this equation. First, re-write the equation as Sqrt(46+1)=3+Sqrt(y-2). Square both sides. Unfortunately, you will still have 1 instance of Sqrt(y-2) in your result. Isolate this Sqrt(y-2), and then square both sides again. Show all work, please.

OpenStudy (legomyego180):

yup youre right

OpenStudy (mathmale):

Yes: I'm with Lord_Box. Again, please show your work..

OpenStudy (mathmale):

Sqrt(4y+1)=??

OpenStudy (republican31):

Sorry I had to move along in the assignment and come back to this. Now to work on it.

OpenStudy (republican31):

Sqrt(4y+1)=3+Sqrt(y-2)

OpenStudy (republican31):

4y+1=9+6Sqrt(y-2)+y-2 ?

TheSmartOne (thesmartone):

Good, now isolate sqrt(y-2)

OpenStudy (republican31):

Ok. 4y+1=7+6Sqrt(y-2)+y 4y-6=6Sqrt(y-2)+y 3y-6=6Sqrt(y-2) and I'm stuck and not even sure if I'm in right direction.

TheSmartOne (thesmartone):

you're correct, now divide 6 on both sides and then you square both sides :)

OpenStudy (republican31):

1/2y-1=Sqrt(y-2)???

TheSmartOne (thesmartone):

yes, now square both sides (:

OpenStudy (republican31):

1/4y+1=y-2 ? This was the route I was on when I failed this question but it didn't quite work out.

TheSmartOne (thesmartone):

(1/2y-1)^2 using (a - b)^2 = a^2 - 2ab + b^2 you made an error when doing it :)

OpenStudy (goldphenoix):

\[\sqrt{4y+1}-\sqrt{y-2}=3\] \[\sqrt{4y+1}=3+\sqrt{y-2}\] \[(\sqrt{4y+1})^2=(3+\sqrt{y-2})^2\] \[4y+1=(3+\sqrt{y-2})^2\] \[4y+1=y+7+6\sqrt{y-2}\] \[3y=6+6\sqrt{y-2}\] \[y=\frac{ 6+6\sqrt{y-2} }{ 3 }\] \[y=2 + 2\sqrt{y-2}\] \[(y-2)^2=(2\sqrt{y-2})^2\] \[y^2-4y+4=4y-8\] \[y^2-8y+12=0\] \[(y-6)(y-2)=0\] \[y= 2,6\] Im pretty sure my method isn't the most quickest, but it gets the job done. Hopefully you understand the steps I took to find the answer

OpenStudy (republican31):

So that means 1/4y^2-y+1? It's just confusing to me. Also thank you GoldPhenoix, that's an intersting method. I'll keep it in mind.

TheSmartOne (thesmartone):

1/4y^2-y+1= y -2 1/4y^2 -2y + 3 = 0 multiply 4 on both sides to get a better coefficient y^2 - 8y + 12 = 0 and we have a quadratic equation to solve :)

OpenStudy (republican31):

Thank you so much all of you!!!! I finally understand! Thank you!!!

TheSmartOne (thesmartone):

Anytime! :)

OpenStudy (republican31):

Now also Sqrt(4y+5)-Sqrt(y-1)=3

OpenStudy (republican31):

Sqrt(4y+5)=3+Sqrt(y-1) 4y+5=9+6Sqrt(y-1)+y-1

TheSmartOne (thesmartone):

yes, now isolate 6sqrt(y-1) and then divide both sides by 6 and then square both sides :)

OpenStudy (republican31):

4y+5=8+6Sqrt(y-1)+y 3y+5=8+6Sqrt(y-1) 3y-3=6Sqrt(y-1)

OpenStudy (republican31):

1/2y-1/2=Sqrt(y-1) 1/4y^2-1/2y-1/4=y-1

OpenStudy (republican31):

y^2-2y-1=y-1 y^2-3y=0 y(y-3)=0 y={0,3} is what I got, but {1,5} is the answer. I don't understand.